OpenStudy (anonymous):

alright. here we go. a long one. verify : sin^5xcos^2x = (cos^2x-2cos^4x+cos^6x)sinx

OpenStudy (psymon):

Turn all but one of the sines into cosines using the pythagorean identity and then multiply it out. \[\sin^{2}x + \cos^{2}x = 1\] \[\sin^{2}x = 1-\cos^{2}x\]

OpenStudy (anonymous):

Mmm.. do i factor anything out first?

OpenStudy (psymon):

Well, you first rewrite it like this: \[\sin^{2}x = 1-\cos^{2}x \implies \sin^{5}xcos^{2}x= sinx(1-\cos^{2}x)^{2}(\cos^{2}x)\]Just dont multiply in that sinx, keep it out. If you only multiply out the cosines you should have it.

OpenStudy (anonymous):

alright, thanks!

OpenStudy (psymon):

Yep, np ^_^